# Rob A. J. Eyles-Ferris,1 Andrew King,1,2,3 Rhaana L. C. Starling,1 Peter G. Jonker,4 Andrew J. Levan,4 Antonio Martin-Carrillo,5 Tanmoy Laskar,6,4 Jillian C. Rastinejad,7,8 Nikhil Sarin,9,10 Nial R. Tanvir,1 Benjamin P. Gompertz,11,12 Nusrin Habeeb,1 Paul T. O’Brien1 and Massimiliano De Pasquale13

### v2

@@ 2.2 Chandra ACIS-S
Figure 1: The long term X-ray light curve of GRB 250702B compared to the prediction of our WD-IMBH TDE model (blue solid line).
@@ 3.2 A wide angle, low Lorentz factor jet
Instead of the ‘frozen-in’ approximation, we use a semi-analytical approximation for the fallback rate which partially takes into account the effects of tidal distortion. To briefly summarise, the frozen-in approximation assumes the specific orbital energy distribution of the debris stream is fixed from the time of disruption. However, numerical simulations have shown this not to be the case due to the tidal distortion of the star at the time of its disruption and the self-gravity of the stream. We therefore use a semi-analytical approximation for the specific energy distribution given by
@@
i.e. three Gaussians with an additional constant for the central region. These represent the ‘wings’ and core of the star. We fit this model to the numerical results of (Cufari22a) using a least squares method and present our results in Figure 2. Figure 2 shows that our semi-analytical approximation accurately recreates the structure of the numerical result and is a significant improvement on the frozen-in approximation. While the masses at the extremes of the orbital specific energy are underestimated, the effect of this is negligible over the timescales we focus on. The more circular WD orbit required here will shift this distribution towards the negative with an offset of $(1-e)/(1-e_{\rm crit}^{-})$ where $e_{\rm crit}^{-}=1-\frac{2r_{t}}{r_{p}}\left(\frac{M_{\bullet}}{M_{\rm WD}}\right)^{-1/3}$ is the critical eccentricity (Hayasaki18). At a given time, $t$, material with a specific orbital energy will return to the IMBH,
@@
Figure 2: Top: the specific orbital energy distribution of a disrupted WD derived using the frozen-in approximation (black), numerical simulations (Cufari22a, red, C22) and our semi-analytical approximation (blue). An eccentricity of $e=1$ is assumed here. Bottom: the ratio between our semi-analytical approximation and the numerical result of (Cufari22a).
We assume the final disruption dominates the long term accretion rate as the earlier stripping episodes are both quicker and will result in significantly less mass being accreted. We fit our model to the long term X-ray decay using the MCMC sampler emcee v3.1.6 (ForemanMackey13) over 5000 iterations with the first 500 discarded as burn-in. In Figure 1, we show 100 traces from the resulting MCMC chain and demonstrate that our model can indeed reproduce the long term evolution of GRB 250702B. We summarise the posterior distributions in Table 1 and also find that our model naturally converges to peak roughly coincident with the third Fermi trigger. However, our model does overpredict the observed luminosity from EP by a factor of two to ten. This is not entirely unexpected, however, and suggests that energy is being lost to some other process rather than being radiated in X-rays. In Section 3.3, we discuss the possibility that this missing energy powers inverse Compton upscattering.
Table 1: The posterior distributions from our MCMC fit to GRB 250702B’s X-ray light curve.
Parameter Posterior
$\log\left(M_{\bullet}/M_{\odot}\right)$ $4.02\pm 0.63$
$M_{\rm WD}$ $0.45\pm 0.20$
$e$ $0.923\pm 0.027$
$\epsilon_{T}$ $0.50\pm 0.25$
We can also use the assumed radiative efficiency of the model to make inferences about the jet properties. We make the conservative assumption that $\epsilon_{X}=0.001$ (Crumley16 find 0.010.10.01-0.1 while including beaming efficiency) and find that $\theta_{j}=0.09^{+0.04}_{-0.02}$ rad. The Lorentz factor is therefore $\Gamma\sim 1/\theta_{j}=11.2^{+2.5}_{-3.3}$. We note that these values are sensitive to our assumed $\epsilon_{X}$ although our assumption is reasonable. Such a jet is wider than typically observed in long GRBs (e.g. Frail01; Yi17), but could be an outlier to the population, and comparable to that inferred for the relativistic TDE Swift J1644+57 at early times (Metzger12; Zauderer13) although we note that Beniamini23 find a significantly larger $\theta_{j}$ for that TDE. It therefore seems feasible that the inferred jet could form in a WD-IMBH TDE. While the Lorentz factor is small, it will still be sufficient to produce a radio counterpart through external shock induced synchrotron although we do not examine it in detail here nor make a direct comparison with the extensive radio observations of GRB 250702B already performed (GCN40979; GCN40985; GCN41053; GCN41054; GCN41059; GCN41061; GCN41145; GCN41147; GCN41215; Levan25). By analogy with other relativistic TDEs (e.g. Pasham15; Hammerstein25), the NIR counterpart will trace thermal emission, possibly from an inflated accretion flow or cooling envelope (e.g. Coughlin14; Dai18; EylesFerris22; Metzger22; Sarin24). However, this is likely to have a peak bolometric luminosity $\lesssim 10^{45}$ erg s-1 and follow the X-ray in fading rapidly, as shown by Levan25. The non-detection of significant transient light by (Gompertz25) is therefore also consistent with this model.
@@ 3.3 Stream induced shocks and gamma-ray emission
The energy released by these collisions is likely to vary somewhat but can be roughly constrained using the formalism of Darbha19999While Darbha19 concentrate on ultra-deep encounters where a single debris stream promptly intersects with itself, the same physics are applicable to similar debris streams interacting with each other.. Near the pericentre, the tidal forces exerted by the black hole will compress the streams (e.g. Luminet85) to a radius $\sim R_{\rm WD}$ and the surface density will therefore be comparable to the original WD i.e. $\rho\sim 3M_{\rm WD}/4\pi R_{\rm WD}^{3}$. The kinetic energy of the shocked material is then given by $\rho v_{s}^{2}$ where $v_{s}$ is the shock velocity which is comparable to the pericentric velocity $v_{s}\sim v_{p}=\left(GM_{\bullet}/r_{p}\right)^{1/2}$. We assume that $r_{p}=r_{t}$ and therefore for our derived posteriors, $v_{p}=0.23^{+0.21}_{-0.12}~c$. The energy is deposited into an optically thin surface layer with thickness $\Delta R\lesssim(cR_{\rm WD}/\rho\kappa v_{p})^{1/2}$, where $\kappa$ is the opacity, and therefore the total energy released thermally is given by
@@
where $l_{c}\sim R_{\rm WD}$. Assuming a standard opacity of $\kappa=0.4$ cm2 g-1, we therefore estimate a luminosity of $4.3^{+9.5}_{-3.2}\times 10^{47}$ erg s-1 emitted over a time $\sim 0.03$ s derived from the light crossing time of the $R_{\rm WD}$. Assuming a perfect blackbody, the spectrum will peak in the medium to hard X-rays at $15.9^{+7.7}_{-6.0}$ keV.
These collisions may not make a significant contribution to the observed X-ray light curve and despite the relatively high peak energy, are still significantly below the spectrum observed by Fermi. They are too faint by over an order of magnitude to trigger Swift’s Burst Alert Telescope assuming a 15 - 150 keV sensitivity of $3\times 10^{-9}$ erg cm-2 s-1 (e.g. Lien16). However, in addition to being observed directly, the emission from the collisions provides a significant field of seed photons that could be upscattered by the jet [^10] The accretion disk and any disk winds may also provide additional sources of seed photons. However, the luminosities of these components and the energies of the resulting photons are unlikely to make a significant contribution. The jet itself may also provide seed photons for synchrotron self-Compton but due to the low peak energy and small $\Gamma$ this may only be a small contribution at gamma-ray energies. We therefore do not examine them further here.. If we make the standard approximation of $\nu\sim\Gamma^{2}\nu_{0}$ (e.g. Rybicki_and_Lightman; Longair), the peak in the inverse Compton spectrum should lie at $2.0^{+2.4}_{-1.4}$ MeV, broadly consistent with the $E_{p}$ identified in the spectral fits of Neights25. This could be due to us underestimating $\Gamma$ by a small factor ($\lesssim 2$). The nature of the seed photon field produced by these shocks can also help explain the highly variable nature of the gamma-rays. As mentioned above, each thermal flash from these collisions will last $\sim 0.03$ s and the chaotic nature of the debris streams will therefore cause significant short timescale variability. It is also worth noting that the emission region of these flashes is at $\sim r_{t}$. For a $M_{\rm WD}=0.45~M_{\odot}$, $M_{\bullet}=10^{4}~M_{\odot}$ system that we assume, this is roughly a lightsecond away from the jet. Variability on timescales of a second in the rest frame is therefore entirely reasonable to expect in the gamma-ray light curve.
@@ 4 Conclusions
A thorough rate calculation is beyond the scope of this Letter, however, the expected rate is broadly compatible with that implied by the uniqueness of this event. While there is no conclusive proof that GRB 250702B was indeed a WD-IMBH TDE and a collapsar remains eminently feasible, we have shown that it is possible within the data currently available to us. Future observations and analysis will help determine the exact properties of the jet and allow us to fully evaluate the inferences we have made here.
@@ Acknowledgements
The data used in this work are publicly available, either from the UKSSDC or the Chandra Data Archive (see footnotes above for links), or Li25.

### v3

@@ 2.2 Chandra ACIS-S
Figure 1: The long term X-ray light curve of GRB 250702B compared to 480 randomly selected traces from our MCMC fit using a WD-IMBH model (blue solid lines). The vertical grey lines mark the times of the Fermi triggers. Note the first plotted WXT point is excluded from our fitting.
@@ 3.2 A wide angle, low Lorentz factor jet
Instead of the ‘frozen-in’ approximation, we use a semi-analytical approximation for the fallback rate which partially takes into account the effects of tidal distortion. To briefly summarise, the frozen-in approximation assumes the differential mass per unit specific orbital energy distribution of the debris stream is fixed from the time of disruption. However, numerical simulations have shown this not to be the case due to the tidal distortion of the star at the time of its disruption and the self-gravity of the stream. We therefore use a semi-analytical approximation for the distribution of differential mass per unit specific orbital energy given by
@@
where $A_{x}$, $\varepsilon_{0,x}$ and $\sigma_{x}$ are the amplitudes, centres and widths of three Gaussians and $k_{c}$ is an additional constant added to the central Gaussian. These represent the ‘wings’ and core of the star. We fit this model to the numerical results of (Cufari22a) using a least squares method and present our results in Figure 2. Figure 2 shows that our semi-analytical approximation accurately recreates the structure of the numerical result and is a significant improvement on the frozen-in approximation. While the differential masses at the extremes of the specific orbital energy are underestimated, the effect of this is negligible over the timescales we focus on. The more circular WD orbit required here will shift this distribution towards the negative with an offset of $(1-e)/(1-e_{\rm crit}^{-})$ where $e_{\rm crit}^{-}=1-2\beta\left(\frac{M_{\bullet}}{M_{\rm WD}}\right)^{-1/3}$ is the critical eccentricity (Hayasaki18), $\beta=\frac{r_{t}}{r_{p}}$ is the impact parameter and $r_{p}$ is the pericentric radius. We assume $\beta=1$ throughout. At a given time, $t$, material with a specific orbital energy will return to the IMBH,
@@
Figure 2: Top: the differential mass per unit specific orbital energy distribution of a disrupted WD derived using the frozen-in approximation (black), numerical simulations (Cufari22a, red, C22) and our semi-analytical approximation (blue). An eccentricity of $e=1$ is assumed here. Bottom: the ratio between our semi-analytical approximation and the numerical result of (Cufari22a).
[…]
Table 1: The priors and posterior distributions from our MCMC fit to GRB 250702B’s X-ray light curve.
Parameter Prior Posterior
$\log\left(M_{\bullet}/M_{\odot}\right)$ $2<\log\left(M_{\bullet}/M_{\odot}\right)<5$ $4.02\pm 0.63$
$M_{\rm WD}$ ($M_{\odot}$) $0.2<M_{\rm WD}<1.4$ $0.45\pm 0.20$
$e$ $\leq 1$ $0.923\pm 0.027$
$\epsilon_{T}$ $0<\epsilon_{T}<1$ $0.50\pm 0.25$
We can also use the assumed radiative efficiency of the model to make inferences about the jet properties. We make the assumption that $0.001\leq\epsilon_{X}\leq 0.01$ (Crumley16 find 0.010.10.01-0.1 while including beaming efficiency) and find that $\theta_{j}\sim 0.07-0.40$ rad. The Lorentz factor is therefore $\Gamma\sim 1/\theta_{j}\sim 2.5-13.7$. We note that these values are sensitive to our assumed $\epsilon_{X}$ although our assumption is reasonable. Such a jet is wider than typically observed in long GRBs (e.g. Frail01; Yi17), but could be an outlier to the population, and comparable to that inferred for the relativistic TDE Swift J1644+57 at early times (Metzger12; Zauderer13) although we note that Beniamini23 find a significantly larger $\theta_{j}$ for that TDE. It therefore seems feasible that the inferred jet could form in a WD-IMBH TDE. While the Lorentz factor is small, it will still be sufficient to produce a radio counterpart through external shock induced synchrotron although we do not examine it in detail here nor make a direct comparison with the extensive radio observations of GRB 250702B already performed (GCN40979; GCN40985; GCN41053; GCN41054; GCN41059; GCN41061; GCN41145; GCN41147; GCN41215; Levan25). By analogy with other relativistic TDEs (e.g. Pasham15; Hammerstein25), the NIR counterpart will trace thermal emission, possibly from an inflated accretion flow or cooling envelope (e.g. Coughlin14; Dai18; EylesFerris22; Metzger22; Sarin24). However, this is likely to have a peak bolometric luminosity $\lesssim 10^{45}$ erg s-1 and follow the X-ray in fading rapidly, as shown by Levan25. The non-detection of significant transient light by (Gompertz25) is therefore also consistent with this model.
@@ 3.3 Stream induced shocks and gamma-ray emission
The energy released by these collisions is likely to vary somewhat but can be roughly constrained using the formalism of Darbha19101010While Darbha19 concentrate on ultra-deep encounters where a single debris stream promptly intersects with itself, the same physics are applicable to similar debris streams interacting with each other.. Near the pericentre, the tidal forces exerted by the black hole will compress the streams (e.g. Luminet85) to a radius $\sim R_{\rm WD}$ and the surface density will therefore be comparable to the original WD i.e. $\rho\sim 3M_{\rm WD}/4\pi R_{\rm WD}^{3}$ assuming a constant density within the WD. The kinetic energy of the shocked material is then given by $\rho v_{s}^{2}$ where $v_{s}$ is the shock velocity which is comparable to the pericentric velocity $v_{s}\sim v_{p}=\left(GM_{\bullet}/r_{p}\right)^{1/2}$. We assume that $r_{p}=r_{t}$ and therefore for our derived posteriors, $v_{p}=0.23^{+0.21}_{-0.12}~c$. The energy is deposited into an optically thin surface layer with thickness $\Delta R\lesssim(cR_{\rm WD}/\rho\kappa v_{p})^{1/2}$, where $\kappa$ is the opacity, and therefore the total energy released thermally is given by
@@
where $l_{c}\sim R_{\rm WD}$ is the light crossing time. Assuming a standard opacity of $\kappa=0.4$ cm2 g-1, we therefore estimate a luminosity of $4.3^{+9.5}_{-3.2}\times 10^{47}$ erg s-1 emitted over a time $\sim 0.03$ s derived from $l_{c}$. Assuming a perfect blackbody, the spectrum will peak in the medium to hard X-rays at $15.9^{+7.7}_{-6.0}$ keV.
These collisions may not make a significant contribution to the observed X-ray light curve and despite the relatively high peak energy, are still significantly below the spectrum observed by Fermi. They are too faint by over an order of magnitude to trigger Swift’s Burst Alert Telescope assuming a 15 - 150 keV sensitivity of $3\times 10^{-9}$ erg cm-2 s-1 (e.g. Lien16). However, in addition to being observed directly, the emission from the collisions provides a significant field of seed photons that could be upscattered by the jet [^11] The accretion disk and any disk winds may also provide additional sources of seed photons. However, the luminosities of these components and the energies of the resulting photons are unlikely to make a significant contribution. The jet itself may also provide seed photons for synchrotron self-Compton but due to the low peak energy and small $\Gamma$ this may only be a small contribution at gamma-ray energies. We therefore do not examine them further here.. If we make the standard approximation of $\nu\sim\Gamma^{2}\nu_{0}$ (e.g. Rybicki_and_Lightman; Longair), the peak in the inverse Compton spectrum should lie in the range $\sim 0.6-4.4$ MeV, depending the value assumed for $\epsilon_{X}$. It is therefore broadly consistent with the $E_{p}$ identified in the spectral fits of Neights25. The nature of the seed photon field produced by these shocks can also help explain the highly variable nature of the gamma-rays. As mentioned above, each thermal flash from these collisions will last $\sim 0.03$ s and the chaotic nature of the debris streams will therefore cause significant short timescale variability. It is also worth noting that the emission region of these flashes is at $\sim r_{t}$. For a $M_{\rm WD}=0.45~M_{\odot}$, $M_{\bullet}=10^{4}~M_{\odot}$ system that we assume, this is roughly a lightsecond away from the jet. Variability on timescales of a second in the rest frame is therefore entirely reasonable to expect in the gamma-ray light curve.
@@ 4 Conclusions
A thorough rate calculation is beyond the scope of this Letter, however, the expected rate is broadly compatible with that implied by the uniqueness of this event. While there is no conclusive proof that GRB 250702B was indeed a WD-IMBH TDE and several models remain eminently feasible, we have shown that it is possible within the data currently available to us. Future observations and analysis will help determine the exact properties of the jet and allow us to fully evaluate the inferences we have made here.
@@ Acknowledgements
The data used in this work are publicly available, either from the UKSSDC or the Chandra Data Archive (see footnotes above for links), or from Li25.